one charge of +5.0μC is placed in air at exactly x=0, and a second charge +7.0C at x=100cm. Where can a third be placed so as to experience zero net force due to the other two


1
Expert's answer
2021-04-06T13:53:59-0400

The electric forces on the third charge "q_3" due to charges "q_1" and "q_2" must be balanced:


"F_{13}=F_{23},""\dfrac{kq_1q_3}{d^2}=\dfrac{kq_2q_3}{(r-d)^2},""\dfrac{q_1}{d^2}=\dfrac{q_2}{(r-d)^2},""q_1(r-d)^2=q_2d^2,""\dfrac{q_1}{q_2}=\dfrac{d^2}{(r-d)^2},""\sqrt{\dfrac{q_1}{q_2}}=\dfrac{d}{(r-d)},""d=\dfrac{r\sqrt{\dfrac{q_1}{q_2}}}{(1+\sqrt{\dfrac{q_1}{q_2}})},""d=\dfrac{1.0\ m\cdot\sqrt{\dfrac{5.0\cdot10^{-6}\ C}{7.0\cdot10^{-6}\ C}}}{(1+\sqrt{\dfrac{5.0\cdot10^{-6}\ C}{7.0\cdot10^{-6}\ C}})}=0.46\ m."